Beth has $2.77 in her coin jar - all in pennies (x), nickels (y) and dimes (z). If she has 53 coins in all and four more nickels than dimes, find how many of each type of coin she has. Setup three equations in x, y, and z and solve the equations to find x(pennies), y(nickels), and z(dimes). Note only one equation has y and z only. I do not even know where to begin, once I find the equation I know how to solve them..I just cant find the equaitons
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



She has 53 coins, so:



The value of her pennies is cents, because pennies are worth 1 cent each.

The value of her nickels is cents, because nickels are worth 5 cents each.

The value of her dimes is cents, because dimes are worth 10 cents each.

The total value of all her coins is $2.77, but since we have expressed all of the values of each type of coins in cents, we should express the total value of her coins in cents as well. $2.77 is 277 cents.

And then:



Finally, we know that there are 4 more nickels than dimes, so



There are your three equations.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

Beth has $2.77 in her coin jar - all in pennies (p), nickels (n) and dimes (d). If she has 53 coins in all and four more nickels than dimes, find how many of each type of coin she has.
-------
Equations:
Coins equation: p + n + d = 53
Value equation: p +5n+10d = 277 cents
Quantity Equation: n-d = 4
======
p + n + d + 53
p +4n+10d = 277
0 + n - d = 4
===================
Solve the system for p,n, and d
----
Note: Your instructions do not give whole number solutions.
Cheers,
Stan H.
============